A consultant charges $45 for each hour she works on a
consultation, plus a flat $30 consulting fee. How many
hours of work are included in a $210 bill for a
consultation?

Answer by mananth(16946)

$45 for each hour
a flat $30 consulting fee.
$210 bill
..
let the number of hours she worked be x
$45 /hour
charges for hours= 45x
..
45x + 30 = 210
add -30 to both sides
45x +30-30=210-30
45x=180
x=180/45
x= 4 the number of hours she works.